hi, need help
i need to find the intercepts of the equation. then graph the equation.
x - 2y = 6 can someone assist me please.
thanks

Answer by wuwei96815(245)

First let y=0:
x-2y=6
x-2(0)=6
x-0=6
x=6

Second let x=0:
x-2y=6
(0)-2y=6
-2y=6
y=6/-2=-3

Therefore, your two coordinates are (6,0) and (0,-3). Now you can plot your line by drawing a straight line through these two points on the graph.